OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. How many pounds of jelly beans and chocolate drops did Terri buy?

OpenStudy (anonymous):

Terri what the hell you can't buy 2 pounds of candy and keep it all to yourself

OpenStudy (anonymous):

i need a real answer please this is timed .

terenzreignz (terenzreignz):

@BeckieBird I think you'll find I can ;)

OpenStudy (anonymous):

hmm ok so if its 2.1 pounds and it cost well she bought at least 1 pound of each,

OpenStudy (anonymous):

yea I figured that out after I posted o.o thanks (::

terenzreignz (terenzreignz):

Really? What's your answer?

OpenStudy (anonymous):

I didnt solve it all the way but I think it was 1/5 pounds of choco drops and the resy jellybeans

OpenStudy (anonymous):

answers could be: 0.31p + 1.41 = 1.56, 0.31p + 1.41 = 2.1, –0.31p + 2.06 = 1.56, –0.31p + 2.06 = 2.1

terenzreignz (terenzreignz):

Okay, we need to formulate the problem... it looks like p is an amount.

terenzreignz (terenzreignz):

If Terri bought p pounds of Jelly Beans, how many pounds of Chocolate drops did she buy?

OpenStudy (anonymous):

here is our possible answers 1.6 pounds of jelly beans; 0.5 pounds of chocolate drops 1.1 pounds of jelly beans; 1 pound of chocolate drops 1 pound of jelly beans; 1.1 pounds of chocolate drops 0.5 pounds of jelly beans; 1.6 pounds of chocolate drops

terenzreignz (terenzreignz):

Fine... the answer is there. Anyway, I asked you a question?

OpenStudy (anonymous):

hold on start over

terenzreignz (terenzreignz):

Okay... Let's assume she bought p pounds of Jelly Beans. Then how many pounds of chocolate drops did she buy?

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. How many pounds of jelly beans and chocolate drops did Terri buy? 1.6 pounds of jelly beans; 0.5 pounds of chocolate drops 1.1 pounds of jelly beans; 1 pound of chocolate drops 1 pound of jelly beans; 1.1 pounds of chocolate drops 0.5 pounds of jelly beans; 1.6 pounds of chocolate drop

terenzreignz (terenzreignz):

You already posted those -_- We're trying to solve it now, but you need to work with me. Now for the third time, if Terri bought p pounds of Jelly Beans, how many pounds of Chocolate drops did she buy?

OpenStudy (anonymous):

i dont know i think she bought 1.6 pounds of chocolte drops

terenzreignz (terenzreignz):

Let's find out together, then?

OpenStudy (anonymous):

ok we need to find out how much 1.6 drops coast for the drops

terenzreignz (terenzreignz):

How did you get 1.6? It's wrong -_-

OpenStudy (anonymous):

1.6 plus 0.5

OpenStudy (anonymous):

1.6 is thwe amount of drops

OpenStudy (anonymous):

0.5 is the jelly beans

terenzreignz (terenzreignz):

how did you arrive at that?

OpenStudy (anonymous):

you have to add the coast of them at that ratio to get the answer

terenzreignz (terenzreignz):

well then.. what did you need me for? -_-

OpenStudy (anonymous):

to check that answer ok

terenzreignz (terenzreignz):

I don't know... -_- It looks iffy.

terenzreignz (terenzreignz):

Anyway, you still have that first question unresolved.